2713:

Marcy is making a hooked circle-shaped rug for her mother. The circumference of the circle-shaped rug is 69.08 inches. What is the area of the rug? Use pi = 3.14.

kittybasil:

@2713 Is there any extra information that would allow us to help you solve any problems? Right now, there is no question to answer.

Vocaloid:

Question has been edited to provide missing info.

darkknight:

So what you have is the circumference. And you need to find the area. To do this, just some conversion needed. The circumference in terms of a radius is \[circumference = 2\pi r\] Once you get r, you can plug that into this formula:\ \[\pi r^2 = area \] Which will give you the area of the rug. Since you use 3.14 for pi, then sub that in into your formulas